# Evaluating Cooling Center Coverage Using Persistent Homology of a Filtered Witness Complex

### v1

@@ Acknowledgments
We thank Mason Porter and Nicole Sanderson for helpful comments and discussions. EO would like to acknowledge the support and funding provided by UCLA’s Queen’s Road Foundation. Lastly, we thank the developers of the GeoPandas [23], Shapely [17], networkx [18], and GUDHI packages [30] for their open source code used for the creation of our maps, analysis, and figures.
@@ Appendix A Data Collection
Below describes the data sources along with any pre-processing steps applied to them. A.1 describes the landmark (census block centroids) data and A.1 describes the witness (cooling center) data. Figure 9 plots the full dataset of landmarks and witnesses.
Figure 9: Final dataset of census tract centroids (landmarks) and cooling centers (witnesses) for each of the four locations, colored in light green and dark green, respectively.
A.1 Landmark Data Collection and Processing
The Shapefiles for mapping the census blocks were retrieved from Data.gov [33]. The following datasets were downloaded in July 2024:
  1. 1.
@@ A.1 Landmark Data Collection and Processing
The dataset of census block centroids were created by using the .centroid attribute of the Shapely library in Python for each of the .json files and plotted in Figure 9 in light green.
A.2 Witness Data Collection and Processing
@@
Table 1: Furthest Southwest and Northeast coordinates used to search OpenStreetMap for cooling center tags.
@@
A.3 Heat Vulnerability Index Data Collection and Processing
The following Shapefiles were downloaded in Feb 2024 to create the HVI score maps:
  1. 1.
@@ A.3 Heat Vulnerability Index Data Collection and Processing
Data on the typical afternoon temperature, the percentage of the area covered in tree canopy, number of individuals who are under 5, and number of individuals who are older than 65 was retrieved from the National Integrated Heat Health Information System (NIHHIS). The dataset was created as part of the National Oceanic and Atmospheric Administration’s (NOAA’s) “U.S. Urban Heat Island Mapping Campaign.” Data on the area not covered in tree canopy was found by subtracting the data on the percent of area covered in tree canopy in a given census tract from 100. We verified the independence of these four variables using variance inflation factor (VIF). Those results are presented in Table 2. For further information on the data variables and the data’s geographical scope refer to [26].
@@
Table 2: Variance inflation factor (VIF) results to check for multicollinearity in the four cities of interst. Note: a VIF of 1 indicates no multicollinearity. VIFs between 1 and 5 are considered to have negligible to minimal multicollinearity.
Variable Austin VIF Boston VIF Miami VIF Portland VIF
Typical PM temp. 1.1452 2.5749 1.2388 4.9190
Tree canopy % 1.0786 2.6938 1.0098 4.8477
Population age $<$ 5 1.2158 1.4188 1.2230 1.3257
Population age $>$ 65 1.2612 1.6367 1.3503 1.3527

### v2

@@ Acknowledgments
We thank Mason Porter and Nicole Sanderson for helpful comments and discussions. We would also like to thank the two anonymous reviewers for their constructive comments and suggestions, which helped improve the paper. EO would like to acknowledge the support and funding provided by UCLA’s Queen’s Road Foundation. Lastly, we thank the developers of the GeoPandas [27], Shapely [20], networkx [21], and GUDHI packages [35] for their open source code used for the creation of our maps, analysis, and figures.
@@ Appendix A Data Collection
In this section we describe the data sources along with any pre-processing steps applied to them. A.1 describes the landmark (census block centroids) data and A.2 describes the witness (cooling center) data. Figure 11 plots the full dataset of landmarks and witnesses.
Figure 11. Final dataset of census tract centroids (landmarks) and cooling centers (witnesses) for each of the four locations, colored in light green and dark green, respectively.
A.1. Landmark Data Collection and Processing
Our approach requires a Shapefile containing some set of standardized geographic units that subdivide the region of interest into discrete, mappable areas. There are many different scales of these geographic units. Three examples of regions based on data collected by the US Census Bureau [39] include:
  • Census blocks: the smallest geographic division for which the US Census Bureau collects data. The population of each block is approximately equivalent to that of a city block in an urban area.
  • Census block groups: groups of contiguous census blocks containing a total of 600–3,000 people.
  • Census tracts: groups of contiguous census block groups containing a total of 2,500-8,000 residents.
For our analysis using persistent homology, we use census blocks as they are allow for the most granular level of detail. However, one could use larger geographic units such as block groups or tracts.
The Shapefiles to map the census blocks were retrieved from Data.gov [40]. The following datasets were downloaded in July 2024:
  1. (1)
@@ A.1. Landmark Data Collection and Processing
From the Shapefile, we compute the centroid (the geographic center) of each census block using the .centroid attribute of the Shapely library in Python.
A.2. Witness Data Collection and Processing
@@
Table 3. Furthest Southwest and Northeast coordinates used to search OpenStreetMap for cooling center tags.
@@
A.3. Heat Vulnerability Index Data Collection and Processing
The following Shapefiles were downloaded in February 2024 to create the HVI maps:
  1. (1)
@@ A.3. Heat Vulnerability Index Data Collection and Processing
Data on the typical afternoon temperature, the percentage of the area covered in tree canopy, number of individuals who are under 5, and number of individuals who are older than 65 was retrieved from the National Integrated Heat Health Information System (NIHHIS). The dataset was created as part of the National Oceanic and Atmospheric Administration’s (NOAA’s) “U.S. Urban Heat Island Mapping Campaign.” Data on the area not covered in tree canopy was found by subtracting the data on the percent of area covered in tree canopy in a given census tract from 100. Figures 1215 show the the individual HVI variables for each city.
Our motivation for incorporating these particular four variables into our measure of vulnerability is detailed as follows. Firstly, one can imagine that a warmer census block with few trees may not fare well during an extreme heat event. Secondly, being over the age of 65 puts an individual more at risk of heat related mortality due to social isolation and a lack of mobility to get assistance when faced with emergencies [19]. Lastly, because children thermoregulate less effectively than adults, they are more likely to respond poorly during extreme heat events [19]. The census block is interpreted to be at greater risk when its HVI is higher. See Figure 3 for the maps of the HVI across the four regions of interest. More details on the demographic data used to create the HVI can be found in Appendix A.3.
Appendix B Additional Data Analysis
We tested the independence of the four variables from the HVI using variance inflation factor (VIF). The results are presented in Table 4. These results are discussed in Section 3.2. We further explore the relationships between the HVI variables by considering their pairwise correlations. Figures 1619 show scatterplots of all pairwise relationships and the corresponding As expected, from the VIF analysis, Boston and Portland have a strong pairwise correlation between the typical afternoon temperature and the percentage of the area not covered with tree canopy. In Portland, however, neither variable is strongly correlated with either demographic variables.
Austin
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12. Heatmaps of the four HVI variables for Austin. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Boston
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13. Heatmaps of the four HVI variables for Boston. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Miami
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14. Heatmaps of the four HVI variables for Miami. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
Portland
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15. Heatmaps of the four HVI variables for Portland. Top left: area not covered by tree canopy; top right: typical afternoon temperature; bottom left: number of individuals of age less than 5 years old; bottom right: number of individuals of age greater than 65.
@@ Appendix B Additional Data Analysis
Table 4. Variance inflation factor (VIF) results to check multicollinearity in the four cities of interest. Note: a VIF of 1 indicates no multicollinearity. VIFs between 1 and 3 are considered to have negligible to minimal multicollinearity [22].
Variable Austin Boston Miami Portland
Typical Afternoon Temp. 1.1452 2.5749 1.2388 4.9190
% No Tree Canopy 1.0786 2.6938 1.0098 4.8477
Population age << 5 1.2158 1.4188 1.2230 1.3257
Population age >> 65 1.2612 1.6367 1.3503 1.3527
Table 5. The mean, median, and standard deviation (SD) of the HVI values across all census tracts within a given city.
City Mean Median SD
Austin 7.819 7.427 1.746
Boston 7.326 7.26 1.499
Miami 8.715 8.182 2.051
Portland 7.046 7.076 1.805
Figure 16. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Austin; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.
Figure 17. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Boston; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.
Figure 18. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Miami; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value value of $p<0.05$.
Figure 19. (a) Histograms and pairwise scatter plots of HVI variables across census tracts in Portland; (b) Pairwise Pearson correlations between the specified variables. A star indicates the correlation is not statistically significant as it does not have a $p$-value of $p<0.05$.